# Moving Mesh with Streamline Upwind Petrov-Galerkin (MM-SUPG) Method for Time-dependent Convection-Dominated Convection-Diffusion Problems

### v3

@@ 5 Summary and Conclusions
In this paper, we have developed a moving mesh SUPG (MM-SUPG) method that integrates the moving mesh partial differential equation (MMPDE) methods with streamline upwind Petrov-Galerkin (SUPG) method for convection-diffusion problems. MM-SUPG method takes the advantage of the SUPG method to stabilize the numerical solution and the moving mesh method to adapt the mesh to improve the efficiency of the computation and accuracy of the solution. We considered convection-dominated problems where sharp layers exist in the solution. Both time-dependent and time-independent flows are investigated. And the method is designed to work with both isotropic and anisotropic diffusion tensors.
For isotropic diffusion case, our MM-SUPG method shows a $O(h^{1})$ convergence rate in the $H^{1}$-seminorm and $O(h^{1.5})$ convergence in the $L^{2}$ norm (see Figures 1 and 4). Even with moderately large $N$, we see in Figure 5 that oscillations can still be present when using moving mesh with standard Galerkin finite element method (MM-FEM). However, when using MM-SUPG method, the oscillations are significantly reduced and the solution is much smoother. The results in Examples 3.1 and 3.2 demonstrate the advantage of MM-SUPG over fixed mesh with SUPG (FM-SUPG) and moving mesh without SUPG (MM-FEM).
For anisotropic diffusion case, we have developed conditions for mesh and time step size so that the numerical solution satisfies the discrete maximum principle. Based on the diffusivity tensor $\mathbb{D}$ and the flow vector $\mathbf{b}$, we have introduced a weighted diffusion tensor $\mathbb{K}$ in (50) and developed a new metric tensor, $\mathcal{M}_{int}$ in (115) to control the mesh. Numerical results in Examples 4.1 and 4.2 demonstrate that $\mathcal{M}_{int}$ mesh provides better solution than other meshes in terms of DMP satisfaction.
In summary, the MM-SUPG method takes the advantage of both SUPG method and moving mesh method to provide a stable and accurate solution for convection-diffusion problems. Together with conditions for mesh and time step size, the MM-SUPG method can provide a solution that satisfies the discrete maximum principle, especially when the diffusion tensor is anisotropic.

### v4

@@ 5 Summary and Conclusions
In this paper, we have developed a moving mesh streamline upwind Petrov-Galerkin (MM-SUPG) method that combines the moving mesh partial differential equation (MMPDE) approach with the SUPG stabilization for convection-diffusion problems. The key feature of the proposed method is not only the combination of these two techniques, but also their interaction: the evolving mesh modifies the local element geometry and hence influences both the direction and magnitude of the SUPG stabilization, while the stabilization improves robustness in convection-dominated regimes.
The MM-SUPG method leverages the complementary roles of the two components. The moving mesh method concentrates mesh elements in regions with sharp gradients, improving resolution and computational efficiency, while the SUPG method stabilizes the numerical solution by adding streamline diffusion. The method is applicable to both time-dependent and time-independent problems, and is designed to handle both isotropic and anisotropic diffusion tensors.
For the isotropic diffusion case, the MM-SUPG method exhibits approximately $O(h^{1})$ convergence in the $H^{1}$-seminorm and $O(h^{1.5})$ convergence in the $L^{2}$ norm (see Figures 6 and 7). Numerical experiments show that, for convection-dominated problems with sharp layers, moving mesh with the standard Galerkin method (MM-FEM) may still exhibit noticeable oscillations on moderately refined meshes (see Figure 1). The use of SUPG stabilization within the moving mesh framework reduces these oscillations and leads to improved stability. Comparisons in Examples 3.1 and 3.2 indicate that MM-SUPG improves robustness over fixed mesh SUPG (FM-SUPG), and provides comparable accuracy to MM-FEM in cases where mesh adaptation alone already captures the solution well.
For the anisotropic diffusion case, we have derived sufficient conditions on the mesh geometry and time step size under which the fully discrete scheme satisfies the discrete maximum principle (DMP). The analysis is based on quantitative bounds for the convection terms rather than sign arguments, and therefore avoids restrictive assumptions on the divergence of the velocity field. Based on the diffusivity tensor $\mathbb{D}$ and the flow vector $\boldsymbol{b}$, we introduce a weighted tensor $\mathbb{K}$ in (58), which serves as an analytical tool for incorporating the interaction between convection and diffusion. The tensor $\mathbb{K}$ is not guaranteed to be symmetric positive definite in general; thus, the DMP analysis is established under structural assumptions on $\boldsymbol{b}$ and $\mathbb{D}$ such that its elementwise form is of SPD type.
To further improve monotonicity properties, we construct a metric tensor $\mathcal{M}_{int}$ in (125) through metric intersection, combining information from both $\mathbb{D}$ and $\mathbb{K}$. Numerical results in Examples 4.1 and 4.2 show that meshes generated using $\mathcal{M}_{int}$ improve DMP behavior, as evidenced by reduced undershoots, while differences in standard error measures remain modest.
In summary, the MM-SUPG method provides a stable and effective approach for convection-dominated convection-diffusion problems by combining mesh adaptation and stabilization in a unified framework. Under suitable conditions on the mesh, problem data, and time step size, the method can produce solutions that satisfy the discrete maximum principle, particularly in anisotropic diffusion settings. The results demonstrate that the interaction between mesh adaptation and stabilization plays a key role in achieving both accuracy and improved monotonicity properties.